# Quasi-fuchsian flows and the coupled vortex equations

### v1

@@ Abstract.
We provide an alternative construction of the quasi-Fuchsian flows introduced by Ghys in [Ghy92]. Our approach is based on the coupled vortex equations that allows to see these flows as thermostats on the unit tangent bundle of the Blaschke metric uniquely determined by a conformal class and a holomorphic quadratic differential. We also give formulas for the marked length spectrum of a quasi-Fuchsian flow.
@@ 1. Introduction
If $A\in C^{\infty}(M;\mathcal{K}^{\otimes m})$, we may set $\lambda=\operatorname{Im}(\pi_{m}^{*}A)$, and we are interested in the flow generated by the vector field $F:=X+\lambda V$. The pair $(g,A)$ is said to satisfy the coupled vortex equations if:
@@
such that if we let $([g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}}):=\mathcal{G}([g]_{\mathrm{conf}},A)$, then the weak stable foliation of the hyperbolic metric $g_{1}$ is smoothly conjugate to the foliation defined by (1.2) for $r^{s}$, and the weak unstable foliation of the hyperbolic metric $g_{2}$ is smoothly conjugate to the foliation defined by (1.2) for $r^{u}$. The bijection $\mathcal{G}$ will be constructed using the harmonic map approach to Teichmüller theory (see [EE69, Tro92]). Given two hyperbolic metrics $\sigma$ and $\rho$ on $M$, it is well-known that there exists a unique harmonic diffeomorphism $f:(M,\sigma)\to(M,\rho)$ isotopic to the identity. Moreover, the Hopf quadratic differential $(f^{*}\rho)^{2,0}$ is holomorphic (here the $(2,0)$-part of $f^{*}\rho$ is taken with respect to the complex structure derived from $[\sigma]_{\mathrm{conf}}$), and thus we have a map
@@
where $\text{QD}([\sigma]_{\mathrm{conf}})$ denotes the vector space of holomorphic quadratic differentials on the Riemann surface determined by $[\sigma]_{\mathrm{conf}}$. This map is studied in [Sam78, Wol89], where it is shown that it is a homeomorphism, see [Wol89, Theorem 3.1]. (Note that this map only depends on the conformal class $[\sigma]_{\mathrm{conf}}$, which is being used as a base point; this is another way of showing that ${\mathcal{T}}(M)$ is homeomorphic to $\mathbb{R}^{6g(M)-6}$, where $g(M)$ denotes the genus of $M$.)
@@
Theorem 1.1 provides an independent “PDE proof” of [Ghy92, Théorème B], and the classification result [Ghy93, Théorème 4.6] shows that the two constructions are equivalent up to smooth orbit equivalence. It is interesting to note that Ghys’ proof of Item d relies on the marked length spectrum rigidity for hyperbolic metrics, while in our case we invoke [MP19, Theorem 5.5], which relies on an energy identity and ideas from geometric inverse problems. To show that $\mathcal{G}$ is a bijection, we consider a functional given by the sum of two Dirichlet energies and use the argument in the proof of [Sch93, Proposition 2.12] to establish that this functional has a unique critical point (a global minimum).
@@
Often it is more practical to describe the points $[g_{1}]_{\mathrm{Teich}}$ and $[g_{2}]_{\mathrm{Teich}}$ using the marked length spectrum. As a simple addendum to the main theorem we will show that the $g_{1}$-length of the unique closed geodesic of $g_{1}$ in the non-trivial free homotopy class $\mathfrak{c}$ is given by
@@
where $\zeta$ is the unique closed orbit of $F$ whose projection to $M$ is in the class $\mathfrak{c}$ (see Section 5), and $\ell_{g}(\pi\circ\zeta)$ is the $g$-length of this projection (which is easily seen to agree with the period of $\zeta$). Similarly, the marked length spectrum for $g_{2}$ is given by
@@
It follows right away from (1.4) and (1.5) that the marked length spectrum of $F$ is the arithmetic mean of the marked length spectra of the geodesic flows of $g_{1}$ and $g_{2}$:
$\ell_{g}(\pi\circ\zeta)=\frac{1}{2}(\ell_{g_{1}}(\mathfrak{c})+\ell_{g_{2}}(\mathfrak{c})).$
@@
This paper is organized as follows: Section 2 provides preliminaries on the geometry of the unit circle bundle, the coupled vortex equations, harmonic maps, and the Hopf differential. Section 3 explains why the map $\mathcal{G}$ is a bijection. Section 4 introduces the explicit conjugacy between foliations and completes the proof of Theorem 1.1. The final Section 5 provides proofs of the marked length spectrum identities (1.4) and (1.5).
@@ 3. The bijection $\mathcal{G}$
By (2.11), $\mathcal{G}$ is invariant under the $\text{Diff}_{0}(M)$-action and hence it descends to a map on $\mathcal{T}\mathrm{QD}(M)$ (still denoted by $\mathcal{G}$). The next proposition paves the way for the proof of Theorem 1.1.
@@
Let $(M,[\sigma]_{\mathrm{conf}})$ be a Riemann surface of genus $\geq 2$ and $g_{1}$ a hyperbolic metric. As before, we let $f:(M,[\sigma]_{\mathrm{conf}})\to(M,g_{1})$ be the unique harmonic diffeomorphism isotopic to the identity and $A:=(f^{*}g_{1})^{2,0}$ the holomorphic Hopf differential associated to $f$. Recall that (see (2.3)) we denote by $g$ the Blaschke metric determined by the data $([\sigma]_{\mathrm{conf}},A)$ so that $g$ and $A$ are linked by $K_{g}=-1+|A|^{2}_{g}$. We start with the following observation which is probably well known to experts, compare with [Hit87, Theorem 11.2]; we include a proof for completeness.
@@
But for the hyperbolic metric $h$, the weak stable bundle is given by $\ker(\beta_{1}+\psi_{1})$ and the claim follows; the constructed conjugacy is explicitly $df\circ\mathcal{I}:SM_{g}\to SM_{g_{1}}$. Running the same construction with $-A$ and $g_{2}$ (which amounts to switching $\lambda$ by $-\lambda$) we obtain a conjugacy between weak unstable foliations thus proving Items a, b, and c in Theorem 1.1. The proof of Item d is a direct consequence of [MP19, Theorem 5.5]. ∎
@@
5.2. Equality of marked length spectra

### v2

@@ Abstract.
We provide an alternative construction of the quasi-Fuchsian flows introduced by Ghys in [Ghy92]. Our approach is based on the coupled vortex equations that allows to see these flows as thermostats on the unit tangent bundle of the Blaschke metric uniquely determined by a conformal class and a holomorphic quadratic differential. We also give formulas for the marked length spectrum of a quasi-Fuchsian flow in the thermostat parametrization.
@@ 1. Introduction
If $A\in C^{\infty}(M;\mathcal{K}^{\otimes m})$, we are interested in the flow generated by the vector field
(1.1) $F:=X+\lambda V,\quad\lambda=\operatorname{Im}(\pi_{m}^{*}A).$
The pair $(g,A)$ is said to satisfy the coupled vortex equations if:
@@
such that if we let $([g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}}):=\mathcal{G}([g]_{\mathrm{conf}},A)$, then the weak stable foliation of the hyperbolic metric $g_{1}$ is smoothly conjugate to the foliation defined by (1.3) for $r^{s}$, and the weak unstable foliation of the hyperbolic metric $g_{2}$ is smoothly conjugate to the foliation defined by (1.3) for $r^{u}$. The bijection $\mathcal{G}$ that we will use is well known in the literature from different points of view. In this note we choose the harmonic map approach to Teichmüller theory (see [EE69, Tro92]) but we could have also used the Anti-de Sitter (AdS) point of view as originated in the paper by Mess [Mes07]; we briefly survey this alternative approach in Remark 4.2 below.
Given two hyperbolic metrics $\sigma$ and $\rho$ on $M$, it is well-known that there exists a unique harmonic diffeomorphism $f:(M,\sigma)\to(M,\rho)$ isotopic to the identity. Moreover, the Hopf quadratic differential $(f^{*}\rho)^{2,0}$ is holomorphic (here the $(2,0)$-part of $f^{*}\rho$ is taken with respect to the complex structure derived from $[\sigma]_{\mathrm{conf}}$), and thus we have a map
@@
where $\text{QD}([\sigma]_{\mathrm{conf}})$ denotes the vector space of holomorphic quadratic differentials on the Riemann surface determined by $[\sigma]_{\mathrm{conf}}$. This map is studied in [Sam78, Wol89], where it is shown that it is a homeomorphism, see [Wol89, Theorem 3.1]. (Note that this map only depends on the conformal class $[\sigma]_{\mathrm{conf}}$, which is being used as a base point.)
Let us introduce some notation concerning the marked length spectrum. For a non trivial free homotopy class $\mathfrak{c}$ in $M$ and a hyperbolic metric $\sigma$, write $\ell_{\sigma}(\mathfrak{c})$ for the length (with respect to $\sigma$) of the unique closed geodesic in the class $\mathfrak{c}$.
@@
  5. e.
    For any non trivial free homotopy class $\mathfrak{c}$ in $M$, there is a unique closed orbit $\zeta$ of the flow generated by (1.1) such that the projection $\pi\circ\zeta$ of $\zeta$ to $M$ belongs to $\mathfrak{c}$, and
    (1.5) $\ell_{g}(\pi\circ\zeta)=\frac{1}{2}(\ell_{g_{1}}(\mathfrak{c})+\ell_{g_{2}}(\mathfrak{c})),$
    where $\ell_{g}(\pi\circ\zeta)$ is the length of $\pi\circ\zeta$ with respect to $g$.
Theorem 1.1, Items a–d, provide an independent “PDE proof” of [Ghy92, Théorème B], and the classification result [Ghy93, Théorème 4.6] shows that the two constructions are equivalent up to smooth orbit equivalence. We note that the appealing formula in Item e, expressing the marked length spectrum of the thermostat flow as the arithmetic mean of those of $g_{1}$ and $g_{2}$, depends on the particular parametrization of the flow and is a novel feature not detected in [Ghy92] (note that $\ell_{g}(\pi\circ\zeta)$ coincides with the period of $\zeta$). It is interesting to note that Ghys’ proof of Item d relies on the marked length spectrum rigidity for hyperbolic metrics, while in our case we invoke [MP19, Theorem 5.5], which relies on an energy identity and ideas from geometric inverse problems. One possible approach to show that $\mathcal{G}$ is a bijection, is to consider a functional given by the sum of two Dirichlet energies and use the argument in the proof of [Sch93, Proposition 2.12] to establish that this functional has a unique critical point (a global minimum).
@@
Often it is more practical to describe the points $[g_{1}]_{\mathrm{Teich}}$ and $[g_{2}]_{\mathrm{Teich}}$ using the marked length spectrum. In the course of proving the main theorem, we will also show that for a free homotopy class $\mathfrak{c}$ in $M$ we have
@@
where $\zeta$ is as in Item e. Similarly, the marked length spectrum for $g_{2}$ is given by
@@
Item e from Theorem 1.1 follows right away from (1.6) and (1.7).
@@
This paper is organized as follows: Section 2 provides preliminaries on the geometry of the unit circle bundle, the coupled vortex equations, harmonic maps, and the Hopf differential. Section 3 explains why the map $\mathcal{G}$ is a bijection. Section 4 introduces the explicit conjugacy between foliations and proves Theorem 1.1. The final Section 5 provides proofs of the marked length spectrum identities (1.6) and (1.7).
@@ 3. The bijection $\mathcal{G}$
By (2.11), $\mathcal{G}$ is invariant under the $\text{Diff}_{0}(M)$-action and hence it descends to a map on $\mathcal{T}\mathrm{QD}(M)$ (still denoted by $\mathcal{G}$). The next proposition paves the way for the proof of Theorem 1.1; we include a proof for completeness sake.
@@
Let $(M,[\sigma]_{\mathrm{conf}})$ be a Riemann surface of genus $\geq 2$ and $g_{1}$ a hyperbolic metric. As before, we let $f:(M,[\sigma]_{\mathrm{conf}})\to(M,g_{1})$ be the unique harmonic diffeomorphism isotopic to the identity and $A:=(f^{*}g_{1})^{2,0}$ the holomorphic Hopf differential associated to $f$. Recall that (see (2.3)) we denote by $g$ the Blaschke metric determined by the data $([\sigma]_{\mathrm{conf}},A)$ so that $g$ and $A$ are linked by $K_{g}=-1+|A|^{2}_{g}$. We start with the following observation which is well known to experts, compare with [Hit87, Theorem 11.2]; we include a proof for completeness.
@@
Remark 4.2.
It is possible to recast the bijection $\mathcal{G}$ and Lemma 4.1 using AdS geometry; this originates in the work of Mess in [Mes07]. A pair $([g_{1}]_{\mathrm{Teich}},[g_{2}]_{\mathrm{Teich}})$ uniquely determines a maximal global hyperbolic AdS structure on $M\times\mathbb{R}$. It turns out that such an AdS manifold contains a unique Cauchy surface that is maximal, that is, it has zero mean curvature [BBZ07]. In addition its first fundamental form matches precisely the Blaschke metric $g$; one may see this by checking that the coupled vortex equations agree with the Gauss-Codazzi equations for the maximal surface. Moreover, the maximal surface comes equipped with the two “projection” maps that are the unique harmonic diffeomorphisms isotopic to the identity that we described earlier having opposite Hopf differentials. This approach also delivers Lemma 4.1 as explained in [BS20, Section 7]. The bijection $\mathcal{G}$ may also be constructed using AdS geometry, see [KS07, Theorem 3.8] and [BS20, Subsection 7.3].
@@
But for the hyperbolic metric $h$, the weak stable bundle is given by $\ker(\beta_{1}+\psi_{1})$ and the claim follows; the constructed conjugacy is explicitly $df\circ\mathcal{I}:SM_{g}\to SM_{g_{1}}$. Running the same construction with $-A$ and $g_{2}$ (which amounts to switching $\lambda$ by $-\lambda$) we obtain a conjugacy between weak unstable foliations thus proving Items a, b, and c in Theorem 1.1. The proof of Item d is a direct consequence of [MP19, Theorem 5.5]. Finally, Item e is proved in Section 5 below. ∎
Remark 4.6.
The explicit map $\mathcal{I}$ provides an alternative and simpler way to compute the resonant 1-forms at zero for $F$; these were originally computed in [CP25, Theorem 1.5 and Section 8] using vertical Fourier analysis. Indeed, if we let $\text{Res}_{0}^{1}(g_{2})$ denote the space of resonant 1-forms of $g_{2}$ at zero in the kernel of contraction by the geodesic vector field (a well-known object), then the analogous space of resonant $1$-forms $\mathrm{Res}_{0}^{1}(F)$ corresponding to $F$ is given by pull-back by the map $\mathcal{I}$ corresponding to $-A$:
$\text{Res}_{0}^{1}(F)=\mathcal{I}^{*}\text{Res}_{0}^{1}(g_{2}).$
Similarly for the co-resonant 1-forms of $F$ and the metric $g_{1}$. This gives right away that the winding cycles of the two SRB measures of $F$ vanish (something that is not obvious a priori) and that the dimension of resonant (and co-resonant) 1-forms is given by the first Betti number of $M$.
@@
5.2. Equality of marked length spectra and Proof of (1.5)